Outils pour utilisateurs

Outils du site


904_cosmic_febv1

GENERAL INFORMATION

  • AD/DT [ns] - ActiveDelay [ns] / DeadTime [ns];

Parametors: febproto_cern_1; FEB_1112_18; SpillON=6500 SpillOFF=100.

WEBDCS run THR [DACu] HVeff [kV] N trig Veto state Position comment
12027483125007.43YesFEB_1112_18 FEB12 - Internal close to Strip HR
12017483095007.43YesFEB_1112_18 FEB11 - External close to Strip HR
1200748266-748307500Scan: TOP, BOT, BOTH3YesFEB_1112_18 FEB12 - Middle close to Strip center
1194748189-748210495Scan: BOTH2Yes????

Laurent and Christoph went Monday 01/04 ad performed following improvements:

1) Solder the connectors grounding top the strip PCB 2) Connected by tape the LV module to the cassette 3) Added wood below the chamber to avoid contact with table that is not grounded. 4) Realigned

WEBDCS run THR [DACu] HVeff [kV] N trig Veto state Position comment
1205748432-4435036.9-7.92NoFEB_1112_27 FEB12 - Internal close to HR
1206748460-4755086.9-7.92NoFEB_1112_27 FEB12 - Internal close to HR
1207748485-4915136.9-7.92NoFEB_1112_27 FEB12 - Internal close to HR
1208748492-4985106.9-7.92NoFEB_1112_27 FEB12 - Internal close to HR
1213748524-5305136.9-7.92NoFEB_1112_27 FEB12 - Moved 6 cm from internal. Closer to Center by 40 cm.
12145136.9-7.92NoFEB_1112_27 FEB12 - Moved 6 cm from internal. Closer to LR by 80 cm.

5) Added condensators to stabilize the LV (Monday 07)

1216748606-6095136.9-7.82NoFEB_1112_29 FEB12 - Center.
1217748615-6195056.9-7.82NoFEB_1112_29 FEB12 - Center.
1218748620-6265136.9-7.82YesFEB_1112_29 FEB11 - Center.
121974862…5056.9-7.82YesFEB_1112_29 FEB11 - Center.
122074863…5137.43YesFEB_1112_29 FEB11 - Center. High stats, high precision

5) Made shorter the HV cable after the filter (Thursday 11/04)

1222748657-6665136.6-7.82YesFEB_1112_29, feb_proto_2 FEB11 - Center.

6) Added Yuri box using state FEB_1112_30

Lots of scans to be added there

7) Maxime realigned the TDC channels using 6BDac. Spoiled by mistake FEB_1112_30. Recreated state FEB_1112_35.

Unsucsessful attemps: For the state FEB_1112_36 all the even channels piedestals was moved up by 2 6BDac. This shall correspond to 6 DAQu. This shall make LR more sensible.

Second attempt: I used the following mask and built FEB_1112_37

cor12_1= [0 0 0 0 0 2 0 2 0 2 0 2 0 2 0 2 0 2 0 2 0 2 0 2 0 0 2 0 0 0 2 0]

trying to make sense of what I've seen in .xml files

12.1 <DAC6B>0 0 0 0 16 17 20 19 21 19 24 18 21 24 21 23 23 22 20 27 24 25 30 22 29 0 29 0 34 0 31 0 </DAC6B>

and in https://cms.wiki.ipnl.in2p3.fr/debug_febv1_1#mapping_correct_pour_32_cannaux

But when I scan the piedestals I see al channels that moved by 6 DAQu. I'm lost.

MAY LR-collebration

Chamber 90-52 =38 degree from the horizontal line. Middle ~1m of PCB from HR

FEB_1112_39 - all pedestals in same position. FEB_1112_40 - +6 DACu for LR channels.

749172 - thrscan from FEB_1112_39 calibration.

904_cosmic_febv1.txt · Dernière modification: 2019/05/03 12:19 par DAQ user